OpenStudy (anonymous):

What is the equation of the line with an x-intercept of -1 and a y-intercept of 2? y = 2x - 2 y = - 2x + 2 y = 2x + 2

OpenStudy (kidrah69):

y=mx+b m=slope

OpenStudy (anonymous):

i got this

OpenStudy (kidrah69):

m=x so -1 y=2

OpenStudy (kidrah69):

y-intercept

OpenStudy (anonymous):

dang...@kidrah69 stole my answer

OpenStudy (anonymous):

well, i got that much i wasn't sure how to put that information into an equation...

OpenStudy (kidrah69):

Ok y=mx+b right where m is the slope and b is the y intercept

OpenStudy (anonymous):

wooooowww i was thinking way harder than i should of, but also -1 isnt an answer i can pick...

OpenStudy (kidrah69):

There has to be more information than you gave then :/

OpenStudy (kidrah69):

OH

OpenStudy (kidrah69):

Wait but theres no slope :/

OpenStudy (anonymous):

i copied and pasted my question to here..thats all there is..

OpenStudy (anonymous):

is it 2 divided by -1 maybe???

OpenStudy (kidrah69):

No, @amistre64 might know :/

OpenStudy (anonymous):

i have no idea...

OpenStudy (anonymous):

thank you for helping me though!!:) @kidrah69

OpenStudy (amistre64):

we are intercetping the axises; the y intercept is the x=0 axis, let x=0 what is y? the x intercept is the y=0 axis, let y=0, what is x?

OpenStudy (anonymous):

....i dont understand...

OpenStudy (amistre64):

you dont understand which part of it, the addition or the multiplication?

OpenStudy (anonymous):

i didnt understand what you meant for me to do. like how to set up what you were telling me..

OpenStudy (amistre64):

do you have equations to test out?

OpenStudy (anonymous):

my entire question is above... is; y = 2x - 2 y = - 2x + 2 y = 2x + 2

OpenStudy (anonymous):

those are my possible answers

OpenStudy (amistre64):

so they look like: y = mx + b? if so then .... the y intercept is the x=0 axis, let x=0 what is y? the x intercept is the y=0 axis, let y=0, what is x?

OpenStudy (anonymous):

i'm sorry...im really slow when it comes to math...

OpenStudy (amistre64):

do you see an x?

OpenStudy (anonymous):

i see x's in all of the equations

OpenStudy (amistre64):

then turn them into 0s ... what do we end up with?

OpenStudy (anonymous):

ohhhhh i got you!:) just take one of the equations and replace x with 0 find y and do the same thing but switched around??

OpenStudy (amistre64):

yes, not sure what you mean by switch around tho; when you are done finding the ys for x=0 redo them, let y=0 and find the values of x

OpenStudy (anonymous):

yeah thats what i meant like place 0 in for x, find y, then put 0 in for y and find x.

OpenStudy (amistre64):

good, the option that matches the information given in the problem is the correct then one. try it out and we can correct you if need be.

OpenStudy (anonymous):

so when i find x-intercept is -1 and y is 2 thats the right equation. just making sure i fully understand

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thank you so much!:)

OpenStudy (amistre64):

youre welcome

OpenStudy (anonymous):

i got y=2x+2 for an answer